OpenStudy (anonymous):

Quadrilateral ABCD is inscribed in circle N, as shown below. What is the measure of

OpenStudy (jdoe0001):

http://www.mathopenref.com/quadrilateralinscribedangles.html <--- what do you think it may be?

OpenStudy (anonymous):



OpenStudy (anonymous):

@jdoe0001 please help i have 10 mins

OpenStudy (anonymous):

i have ten mins to explain and figure out the answer

OpenStudy (jdoe0001):

well, did you see the "inscribed quadrilateral conjecture"? the opposite angles are 180, or "supplementary angles"

OpenStudy (anonymous):

@jdoe0001 im lost

OpenStudy (jdoe0001):

so... I assume you haven't seen the link above? http://www.mathopenref.com/quadrilateralinscribedangles.html

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

i just don't understand it

OpenStudy (jdoe0001):



OpenStudy (anonymous):

@jhonyy9

OpenStudy (jdoe0001):

solve for "x" then to find angle ABC, just use the found "x" in 4x+1 :)

OpenStudy (anonymous):

how would i solve for x?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

im so lost

OpenStudy (jdoe0001):

hmm well (4x+1) + (2x-7) = 180 remove the parentheses simplify, solve for "x", or "isolate" x

OpenStudy (anonymous):

4+1 2-7?

OpenStudy (anonymous):

5-5?

OpenStudy (jdoe0001):

4+1? seems to assume x =1 but we dunno what "x" is so... solve for "x"

OpenStudy (anonymous):

6x-6=180

OpenStudy (jdoe0001):

yeap 6x - 6 = 180 then 6x = 180 +6 \(x=\cfrac{180+6}{6}\)

OpenStudy (anonymous):

the answer is 125 right?

jhonyy9 (jhonyy9):

how you have got for x 125 ?

jhonyy9 (jhonyy9):

yes right bc x=31 so 4x+1=125 right sure

OpenStudy (anonymous):

nah k you :)

OpenStudy (anonymous):

i can't give u both a medal though